# When Fusion Helps and When It Breaks: View-Aligned Robustness in Same-Source Financial Imaging

### v1

@@ 3.3 View-to-Channel Mapping
For models that take both views as input (denoted as both), we concatenate the two images along the channel dimension. Specifically, we map the OHLCV image view ohlcv to channel 0 (ch0) and the indicator image view indic to channel 1 (ch1). Thus, each sample for a both model is represented as a two-channel input tensor $\mathbf{x}=[v_{\texttt{ohlcv}};v_{\texttt{indic}}]\in\mathbb{R}^{2\times 256\times 256}$.
We also consider single-view baselines that take only one view as input, denoted as ohlcv or indic. In these cases, the input has a single channel $\mathbf{x}\in\mathbb{R}^{1\times 256\times 256}$ and only ch0 exists (i.e., the provided view is always placed in ch0). This convention is important for our robustness evaluation: two-view models allow perturbations to target ch0, ch1, or both channels, whereas single-view models can only be perturbed through ch0. This view–channel mapping allows us to define channel-constrained perturbations that target a specific view in the two-view setting, while keeping comparisons with single-view baselines well-defined.
@@ 3.4 Task and Evaluation Metrics
Prediction target.
For each trading day $t$, we form an input sample by extracting a lookback window of the past $L$ trading days, $\{t-L+1,\dots,t\}$, and predict the next-day directional movement at $(t+1)$. We define the next-day simple return as
@@
where $close_{t}$ denotes the closing price on day $t$ and $\mathbb{I}(\cdot)$ is the indicator function. Thus, the task is a binary classification problem with labels $y_{t}\in\{0,1\}$.
Matthews Correlation Coefficient (MCC).
We use MCC as the primary metric because it summarizes the full confusion matrix and remains informative under class imbalance (Matthews, 1975). MCC can be interpreted as a correlation between predicted and true labels, taking values in [1,1][-1,1].
@@ 4 Threat Model and Evaluation Protocol
We study adversarial robustness under test-time adversarial perturbations (evasion attacks). Given a fixed trained model and an input sample, an adversary perturbs the input image(s) within a small budget to induce misclassification.
@@ 4.1 Threat Model
$\ell_{\infty}$-bounded perturbations (defined in raw pixel space).
Let $\mathbf{x}_{\mathrm{raw}}\in[0,1]^{C\times 256\times 256}$ denote the raw input after scaling 8-bit images to $[0,1]$ (with $C=1$ for single-view and $C=2$ for two-view). The adversary outputs $\mathbf{x}^{\prime}_{\mathrm{raw}}=\mathbf{x}_{\mathrm{raw}}+\boldsymbol{\delta}$ subject to (Szegedy et al., 2014; Goodfellow et al., 2015; Madry et al., 2018)
@@
We evaluate budgets
@@
Optional standardization.
Optionally, the model consumes a standardized input
with $(\boldsymbol{\mu},\boldsymbol{\sigma})$ estimated on the training split. The threat model remains defined in raw space; when attacks are executed in standardized space, we use the equivalent budgets
so that the underlying raw-space constraint in Eq. 3 is preserved. We clip perturbed inputs to the valid range $[0,1]$ (or to $\big[(0-\boldsymbol{\mu})\oslash\boldsymbol{\sigma},\ (1-\boldsymbol{\mu})\oslash\boldsymbol{\sigma}\big]$ when operating in standardized space).
Attacks (untargeted CE).
We consider FGSM (one-step) (Goodfellow et al., 2015) and PGD (multi-step with projection onto the $\ell_{\infty}$ ball) (Madry et al., 2018). Both attacks are untargeted and maximize the cross-entropy loss with respect to the true label, thereby increasing the likelihood of misclassification. FGSM performs a single update followed by clipping to the valid range. PGD uses $K=10$ iterations with step size $\alpha=\epsilon_{\mathrm{adv}}/K$ and projection after each step; we disable random initialization (random_start=False). We note that disabling random start can underestimate worst-case PGD strength; we therefore interpret PGD results in this paper as a consistent, reproducible first-order stress test rather than an exhaustive maximization over random restarts. Full attack hyperparameters are reported in Appendix D.
@@ 4.2 Threat Scenarios and View-Aligned Protocol
Using the view–channel mapping (Section 3.3), we evaluate:
View-constrained perturbations.
For two-view inputs, the adversary perturbs exactly one channel (one view) (Yang et al., 2021):
@@
Joint perturbations.
@@
Fair comparison (view-aligned evaluation).
Single-view models have only one channel and place the provided view in ch0. Therefore, we align comparisons by the attacked view rather than the channel index: (i) ohlcv: single-view ohlcv attacked on ch0 vs. two-view both attacked on ch0; (ii) indic: single-view indic attacked on ch0 vs. two-view both attacked on ch1. Joint perturbations are evaluated only for two-view models.
@@ 4.3 Evaluation and Summary Metrics
Diagnostic metrics for two-view models.
For late-fusion two-view models, we additionally report diagnostic statistics on the clean test set to interpret robustness trends. Specifically, we measure (i) cross-branch agreement and symmetric KL between the two branch predictive distributions, and (ii) branch-wise MCC (computed from each branch alone) versus the fused-head MCC. These diagnostics are used for analysis only and do not affect model selection.
@@ 5 A Stable, Leakage-Resistant Training Baseline
We consider a set of learning and non-learning baselines to evaluate the effectiveness of our data-to-image representations under the same leakage-resistant time-block split protocol. All models perform binary classification for next-day direction prediction and output class logits/probabilities. Depending on the input mode, we use either a single image view (indic or ohlcv) or a two-view input (both) formed by channel-wise stacking after resizing to a common resolution. Unless otherwise stated, all trainable baselines share the same optimization setup, and we select the best checkpoint based on the validation metric.
@@ 5.1 Model Architectures
Compact CNN baseline (Lite-CNN).
Our first baseline is a lightweight convolutional network adapted from the MODII framework (Gao et al., 2022). It follows a shallow CNN design with a small number of convolution and pooling stages to extract local visual patterns from the image representation, followed by a compact MLP classifier head. In the two-view mode (both), we adopt early fusion by channel-wise stacking the two resized views as the input. This model provides a lightweight literature baseline adapted from MODII, enabling a direct comparison under our evaluation protocol.
Pretrained ResNet-18 baseline (ResNet18-P).
To benchmark against a more expressive architecture, we employ a ResNet-18 backbone initialized with ImageNet pretrained weights (He et al., 2016; Deng et al., 2009). The first layer is adapted to support the required number of input channels in each input mode, while the remaining network structure is kept unchanged. In the two-view mode (both), we adopt early fusion by channel-wise stacking the two resized views as the input. This model offers an architecture-level comparison between shallow and deeper CNN backbones under identical settings.
Dual-view consistency variants (*-cons).
In the two-view mode (both), we additionally consider dual-branch late-fusion variants for both backbones, denoted Lite-CNN-Cons and ResNet18-P-Cons. Unlike the early-fusion Lite-CNN/ResNet18-P baselines, *-cons uses dual-branch late fusion. We report three settings for each: $\lambda\in\{0,0.5,1\}$, where $\lambda=0$ disables the consistency term (fused-logit cross-entropy only) and $\lambda>0$ enables cross-view consistency regularization in Sec. 5.2. Unless otherwise stated, we fix $T=2.0$.
Logistic Regression baseline (LogReg).
As a traditional machine-learning baseline, we apply logistic regression on flattened image features. We standardize features and train a linear classifier, which quantifies the extent to which the task can be solved by linear decision boundaries on the constructed representations, without relying on hierarchical spatial feature extraction.
Majority-class (non-learning) baseline (Majority).
We include a non-learning baseline that predicts the single most frequent label computed from the training split. Formally, let $y^{\star}=\arg\max_{c}\#\{i:y_{i}=c,\ (x_{i},y_{i})\in\mathcal{D}_{\text{train}}\}$; the predictor outputs $y^{\star}$ for every input at evaluation time. This baseline has no trainable parameters and serves as a strict lower bound (sanity check) to contextualize gains from learned models.
For reproducibility, implementation details of backbones and fusion heads are provided in Appendix B.
@@ 5.2 Cross-view Consistency Regularization
Motivation.
In the two-view setting (both), each sample provides two aligned image views that describe the same market state (e.g., indic and ohlcv) and share the same next-day direction label. While the two views may emphasize different aspects of the input, their predictions should remain coherent. We therefore introduce a consistency regularizer that encourages agreement between the two view-specific branch predictors.
Dual-branch predictors.
For a two-view input, we employ a dual-branch architecture. Let the branch-specific predictors produce logits $z_{a}$ and $z_{b}$ from the two views, and let a fusion head produce fused logits $z_{f}$ from the concatenated branch features. We use $z_{f}$ as the primary output for the supervised objective, while $z_{a}$ and $z_{b}$ are used to define the consistency term.
Temperature-scaled symmetric KL.
We define temperature-scaled predictive distributions (Hinton et al., 2015; Zhang et al., 2018)
@@
where $T>0$ is a temperature parameter. Our consistency loss is a symmetric KL divergence:
$\mathcal{L}_{\mathrm{cons}}(z_{a},z_{b})=\frac{T^{2}}{2}\Big[\mathrm{KL}(p_{a}\,\|\,p_{b})+\mathrm{KL}(p_{b}\,\|\,p_{a})\Big],$ (11)
where $p_{a}=\mathrm{softmax}(z_{a}/T)$ and $p_{b}=\mathrm{softmax}(z_{b}/T)$.
Training objective.
For the dual-branch two-view model, given the ground-truth label $y\in\{0,1\}$, we optimize
@@
where $\mathcal{L}_{\mathrm{CE}}$ is the standard cross-entropy loss on the fused logits and $\lambda\geq 0$ controls the strength of consistency regularization.
Implementation notes.
We compute the KL terms using numerically stable log_softmax and softmax operations. To verify that the regularizer indeed changes cross-view predictions, we monitor cross-branch agreement and symmetric KL on the clean test set. We also report branch-wise MCC to distinguish whether gains (or failures) come from stronger branches or from the fusion head.
@@ 5.3 Stabilizing Labels with a Minimum-Movement Threshold
Next-day direction labels can be noisy when the realized price change is extremely small. To explicitly study how label ambiguity affects fusion and robustness, we optionally construct ex-post evaluation subsets using a minimum-movement threshold $\tau$. Specifically, for each sample ending at day $t$, we compute the realized next-day return $r_{t+1}$ as in Eq. 1 and discard the sample if $|r_{t+1}|<\tau$. We evaluate $\tau\in\{0,\,0.002,\,0.004,\,0.006,\,0.008,\,0.010\}$ (denoted as min_move).
Protocol note (label-dependent selection).
Because inclusion depends on the realized $|r_{t+1}|$, min_move is a label-dependent sample-selection step and therefore changes the task from unconditional next-day direction prediction to direction prediction conditional on $|r_{t+1}|\geq\tau$. This filtering cannot be applied as an inference-time rule and should not be interpreted as a deployable trading decision. We use it solely to define controlled label-noise regimes for offline benchmarking of fusion behavior and adversarial robustness. To ensure comparability across split variants, filtering is performed before splitting, and all split variants operate on the same subset definition.
@@ 5.4 Leakage-Resistant Time Split with Embargo
Since our inputs are constructed from overlapping sliding windows, proportional random splitting can introduce look-ahead leakage and overly optimistic evaluation; we therefore adopt a leakage-resistant time-block split with embargo for all models (López de Prado, 2018). Trading days are ordered in time and grouped into consecutive blocks of $B$ trading days (block_size); the earliest blocks form the training set, followed by validation blocks, and the latest blocks form the test set. Unless stated otherwise, we use a 70/15/15 split by blocks for train/validation/test.
To mitigate boundary leakage caused by overlapping windows across split boundaries, we apply an embargo of at least $D_{\mathrm{emb}}$ trading days (embargo_days) around boundaries (López de Prado, 2018; Racine, 2000). In practice, we enforce the embargo at the block level by dropping $K=\lceil D_{\mathrm{emb}}/B\rceil$ adjacent blocks to create a temporal gap of at least $D_{\mathrm{emb}}$ days (possibly larger due to block granularity). To keep the test period as intact as possible, we apply a symmetric embargo at the train/validation boundary (dropping $K$ blocks from each side) and primarily drop validation-side blocks at the validation/test boundary. Table 1 summarizes the resulting sample sizes after alignment and min_move filtering, as well as the additional sample reduction induced by the embargo under the same split configuration (block_size=30, embargo_days=20). We report the corresponding split date ranges in the appendix.
min_move Aligned Train/Val/Test Used Embargo Loss
0.000 4842 3390/660/702 4752 1.86%
0.002 3666 2580/480/516 3576 2.45%
0.004 2696 1890/330/386 2606 3.34%
0.006 1989 1380/240/279 1899 4.52%
0.008 1471 1050/150/181 1381 6.12%
0.010 1054 750/90/124 964 8.54%
@@ 5.5 Implementation Details
[…]
Data loading and normalization.
All images are resized to $256\times 256$ and loaded as tensors in $[0,1]$. When enabled (use_norm), we apply per-channel standardization using $(\boldsymbol{\mu},\boldsymbol{\sigma})$ estimated from the training split only, avoiding information leakage from validation/test. Two-view inputs follow the view-to-channel mapping in Sec. 3.3.
Training protocol and leakage control.
We apply min_move filtering (as an offline subset definition) before splitting (Sec. 5.3) and use the leakage-resistant time-block split with embargo (Sec. 5.4), keeping the test time range fixed across experimental comparisons (keep_test_intact=True). Unless stated otherwise, we keep the training hyperparameters (e.g., batch size and dropout) fixed across model/configuration comparisons.
Early stopping.
We train for at most 40 epochs with early stopping (patience=8). To ensure stable optimization across backbones, we use a model-dependent minimum training epoch: min_epoch=15 for Lite-CNN and Lite-CNN-cons, and min_epoch=2 for ResNet18-P and ResNet18-P-cons.
Adversarial evaluation.
Adversarial evaluation is performed at test time only. We consider FGSM (one-step) and PGD (untargeted cross-entropy), and for PGD we use $K=10$ steps with $\alpha=\epsilon_{\mathrm{adv}}/K$ and disable random initialization (random_start=False). We evaluate both view-constrained and joint perturbations under the view-aligned protocol in Sec. 4.2.
Reproducibility.
@@ 6.1 Clean Baselines
Setup and evaluation metric.
We compare a non-learning Majority baseline, a linear Logistic Regression (LogReg) baseline, and two CNN backbones (Lite-CNN and ResNet18-P) under three input modes: ohlcv, indic, and the dual-view both. For both, the vanilla CNN baselines use early fusion via channel-wise stacking, whereas *-cons variants use a dual-branch late fusion head and optionally add cross-view consistency regularization weighted by $\lambda$. Throughout the paper, we use Matthews correlation coefficient (MCC) as the primary and decision metric, since it accounts for all four confusion-matrix entries (TP/TN/FP/FN) and penalizes degenerate “single-direction” predictors, avoiding superficially high accuracy that lacks genuine directional discriminative power.
Minimum-movement filtering defines the evaluation regime.
Figure 2 shows that the minimum-movement threshold $\tau$ (Sec. 5.3) largely determines whether the task is learnable on the ex-post filtered subsets defined by min_move. Without filtering or with weak filtering ($\tau\in\{0,0.002,0.004\}$), most deep models hover around zero MCC (sometimes negative), indicating near-chance behavior on the full set. As $\tau$ increases to $0.006$ and $0.008$, performance improves sharply, consistent with restricting evaluation to subsets that exclude near-zero moves and thus reduce direction-label ambiguity. Aggregating across all deep configurations (Lite-CNN/ResNet18-P and *-cons, all input modes), the mean MCC rises from 0.002-0.002 at $\tau=0$ to $0.082$ at $\tau=0.006$ and peaks at $0.092$ at $\tau=0.008$, before dropping to $0.030$ at $\tau=0.010$. The drop at $\tau=0.010$ aligns with the reduced effective sample size (Table 1), which increases variance and destabilizes training. We stress that $\tau$ is used here to define evaluation regimes ex post, not as an inference-time filtering rule.
Lite-CNN: single-view is more robust under weak filtering; dual-view becomes beneficial once labels stabilize.
Late fusion drives the main gain; consistency regularization yields modest, non-monotonic effects.
ResNet18-P: early fusion underperforms; late fusion partially recovers but variance is higher.
LogReg as a diagnostic baseline.
Takeaways.
Across backbones, (i) label stabilization via min_move is a prerequisite for learning in our setting; (ii) in the dual-view setting, late fusion is generally more reliable than early fusion on clean evaluation once labels stabilize; and (iii) consistency regularization can provide modest gains in some mid-noise regimes but is not reliably monotonic and requires careful tuning jointly with $\tau$.
@@
Model (input) $\boldsymbol{\tau=0.004}$ $\boldsymbol{\tau=0.006}$ $\boldsymbol{\tau=0.008}$
Lite-CNN (both, early fusion) $020004\pm 0.016396 0.0$$088767\pm 0.042662$ $\mathbf{0.107206\pm 0.046632}$
Lite-CNN (indic) $\mathbf{0.107292\pm 0.022477}$ $0.079303\pm 0.032678$ $0.100279\pm 0.048322$
Lite-CNN (ohlcv) $0.003236\pm 0.023789$ $109095\pm 0.039023 0.0$$090707\pm 0.028893$
Lite-CNN-Cons (both, $\lambda=0$) $0.023262\pm 0.034602$ $0.127459\pm 0.035946$ $0.100430\pm 0.034439$
Lite-CNN-Cons (both, $\lambda=0.5$) $0.015634\pm 0.052001$ $\mathbf{0.129607\pm 0.026449}$ $0.089973\pm 0.067209$
Lite-CNN-Cons (both, $\lambda=1$) $0.014013\pm 0.034309$ $0.120755\pm 0.029468$ $0.103269\pm 0.031387$
ResNet18-P (both, early fusion) $012168\pm 0.030674 0.0$$17593\pm 0.044072 0.0$$042873\pm 0.061370$
ResNet18-P (indic) $055638\pm 0.047824 0.0$$082039\pm 0.039988$ $0.108023\pm 0.040145$
ResNet18-P (ohlcv) $005656\pm 0.059240 0.0$$018580\pm 0.054761$ $0.109101\pm 0.072830$
ResNet18-P-Cons (both, $\lambda=0.5$) $003653\pm 0.037078 0.0$$67112\pm 0.031388 0.0$$083423\pm 0.052584$
Choice of $\tau$ for robustness evaluation.
All subsequent adversarial evaluations use $\tau=0.006$ as the default operating point. At smaller thresholds ($\tau\leq 0.004$), most configurations remain close to chance in MCC, making robustness comparisons less informative. At larger thresholds (e.g., $\tau=0.010$), the reduced effective sample size can amplify variance across random seeds. Thus, $\tau=0.006$ offers a practical balance between label stabilization and sufficient data for reliable robustness measurement.

### v2

@@ 3.3 View-to-Channel Mapping
For two-view models (both), we concatenate the two grayscale images along the channel dimension,
$\mathbf{x}=[v_{\texttt{ohlcv}};v_{\texttt{indic}}]\in\mathbb{R}^{2\times 256\times 256},$ (1)
[…]
@@ 3.4 Task and Evaluation Metrics
Prediction target
For each trading day $t$, we form an input sample by extracting a lookback window of the past $L$ trading days, $\{t-L+1,\dots,t\}$, and predict the next-day directional movement at $(t+1)$. In all experiments, we set $L=15$. We define the next-day simple return as
@@
where $close_{t}$ denotes the closing price on day $t$ and $\mathbb{I}(\cdot)$ is the indicator function. Thus, the task is a binary classification problem with labels $y_{t}\in\{0,1\}$. When using the ex-post movement threshold min_move ($\tau$), we report performance on the conditional subset $|r_{t+1}|\geq\tau$; see D for the filtering/split protocol.
Matthews Correlation Coefficient (MCC)
We use MCC as the primary metric because it summarizes the full confusion matrix and remains informative under class imbalance [23]. MCC can be interpreted as a correlation between predicted and true labels, taking values in [1,1][-1,1].
@@ 4 Threat Model and Evaluation Protocol
We evaluate test-time, white-box evasion robustness on the rendered image inputs using standard pixel-space $\ell_{\infty}$ perturbations and first-order attacks [18, 1, 2]. Our goal is a view-aligned stress test: (i) perturb exactly one view for two-view models (view-constrained) or (ii) perturb both views jointly, while keeping comparisons to single-view baselines well defined via the view–channel mapping (Section˜3.3).
@@ 4.1 Threat Model
$\ell_{\infty}$-bounded perturbations in raw pixel space
Let $\mathbf{x}_{\mathrm{raw}}\in[0,1]^{C\times 256\times 256}$ denote the raw input (with $C=1$ for single-view and $C=2$ for two-view). An adversary produces $\mathbf{x}^{\prime}_{\mathrm{raw}}=\mathbf{x}_{\mathrm{raw}}+\boldsymbol{\delta}$ subject to [18, 1, 2]
@@
Budgets
We evaluate a small-budget grid
@@
Attacks
We use FGSM and multi-step PGD as untargeted cross-entropy attacks [1, 2]. All attack hyperparameters (steps, step size, projection, and clipping) are reported in F.
Standardization and valid-range handling
The threat model is always defined in raw $[0,1]$ space. When models consume standardized inputs, we execute attacks in the standardized space using budget conversion that preserves the underlying raw-space constraint, and clip to the corresponding valid range (A).
@@ 4.2 Threat Scenarios and View-Aligned Protocol
Using the view–channel mapping (Section˜3.3), we define two threat scenarios.
View-constrained perturbations (single-view corruption)
For two-view inputs, the adversary perturbs exactly one view/channel [20]:
@@
Joint perturbations (coordinated corruption)
@@
View-aligned comparison to single-view baselines
Single-view models always place the provided view in ch0; therefore, we align comparisons by the attacked view: ohlcv (single-view ohlcv attacked on ch0) vs. (both attacked on ch0), and indic (single-view indic attacked on ch0) vs. (both attacked on ch1). Joint perturbations apply only to two-view models.
@@ 4.3 Evaluation and Summary Metrics
Late-fusion diagnostics (analysis only)
For late-fusion two-view models, we additionally report branch-level diagnostics on the clean test set to aid interpretation; see H.
@@ 5 A Stable, Leakage-Resistant Training Baseline
We consider a set of learning and non-learning baselines to evaluate our image representations under a unified, leakage-resistant evaluation protocol. All models perform binary next-day direction prediction and output logits/probabilities. Inputs are either a single view (indic or ohlcv) or a two-view input (both); for early-fusion models, the two resized views are stacked channel-wise (Section˜3.3). Unless otherwise stated, trainable baselines share the same optimization and checkpoint-selection protocol (E).
@@ 5.1 Model Architectures
Compact CNN baseline (Lite-CNN)
We use a lightweight convolutional network adapted from the MODII framework [11]. In both mode, we use early fusion by channel-wise stacking of the two views; in single-view mode, the model consumes one channel.
Pretrained ResNet-18 baseline (ResNet18-P)
We employ a ResNet-18 backbone initialized with ImageNet pretrained weights [24, 25]. We adapt only the first convolution to match the required number of input channels; the rest of the backbone is unchanged. Two-view both uses early fusion (stacked channels) as above.
Dual-view consistency variants (*-cons)
For two-view inputs, we additionally evaluate late-fusion dual-branch variants for both backbones, denoted Lite-CNN-Cons and ResNet18-P-Cons. These models produce view-specific logits and a fused-logit output; training uses the objective in Section˜5.2. We report $\lambda\in\{0,0.5,1\}$ (with $\lambda=0$ reducing to fused-logit cross-entropy only) and fix $T=2.0$ unless stated otherwise.
Logistic Regression baseline (LogReg)
We apply logistic regression on flattened image features to measure how far linear decision boundaries on the constructed representations can go.
Majority-class baseline (Majority)
We include a non-learning baseline that always predicts the most frequent class in the training split.
For reproducibility, implementation details of backbones and fusion heads are provided in C, including a concise implementation-level summary in Table˜4.
@@ 5.2 Cross-view Consistency Regularization
Motivation
In the two-view setting (both), each sample provides two aligned views of the same market state and shares the same label. We encourage the two view-specific predictors to produce coherent predictions via an explicit consistency regularizer.
Dual-branch setup
Let the two branches produce logits $z_{a}$ and $z_{b}$ from the two views, and let a fusion head produce fused logits $z_{f}$ from concatenated branch features. The fused logits $z_{f}$ define the supervised objective; $z_{a}$ and $z_{b}$ define the consistency term.
Temperature-scaled symmetric KL
We define temperature-scaled predictive distributions [26, 27]
@@
and use the symmetric KL divergence
$\mathcal{L}_{\mathrm{cons}}(z_{a},z_{b})=\frac{T^{2}}{2}\Big[\mathrm{KL}(p_{a}\,\|\,p_{b})+\mathrm{KL}(p_{b}\,\|\,p_{a})\Big].$ (10)
Training objective
We optimize
@@
where $\lambda\geq 0$ controls the regularization strength. Branch-level diagnostics used for interpretation are reported in H.
@@ 5.3 Stabilizing Labels with a Minimum-Movement Threshold
Next-day direction labels can be ambiguous when the realized price change is extremely small. To study how label ambiguity affects fusion and robustness, we optionally define an ex-post evaluation subset using a minimum-movement threshold $\tau$: for each sample ending at day $t$, we compute $r_{t+1}$ (Eq. 2) and keep the sample only if $|r_{t+1}|\geq\tau$. We evaluate $\tau\in\{0,\,0.002,\,0.004,\,0.006,\,0.008,\,0.010\}$ (denoted min_move).
Protocol note
Because inclusion depends on the realized $|r_{t+1}|$, min_move defines a conditional offline benchmark rather than an inference-time decision rule. To keep comparisons controlled across split variants, filtering is applied before splitting so that all variants operate on the same subset definition (D).
@@ 5.4 Leakage-Resistant Time Split with Embargo
Because inputs are built from overlapping sliding windows, random splitting can introduce look-ahead leakage. We therefore adopt a leakage-resistant time-block split with embargo for all models [14]. Trading days are grouped into consecutive blocks of size $B$ (block_size); earlier blocks form train, then validation, and the latest blocks form test (default 70/15/15 by blocks).
To mitigate boundary leakage across split boundaries, we apply an embargo of at least $D_{\mathrm{emb}}$ trading days (embargo_days[14, 28]. Implementation details and the sample-size impact under our default configuration (block_size=30, embargo_days=20) are reported in D (Table˜5).
@@ 5.5 Implementation Details
Training and selection details (optimizer, early stopping, and reproducibility controls) are provided in E. Input standardization and leakage-avoiding normalization are detailed in A, and adversarial-attack hyperparameters are summarized in F.
@@ 6.1 Clean Baselines
Setup and evaluation metric
We compare Majority, LogReg, and two CNN backbones (Lite-CNN and ResNet18-P) under three input modes: ohlcv, indic, and dual-view both. For both, vanilla CNN baselines use early fusion (channel-wise stacking), whereas *-cons variants use a dual-branch late-fusion head and optionally add cross-view consistency regularization weighted by $\lambda$. We use MCC as the primary decision metric since it accounts for all TP/TN/FP/FN entries and penalizes degenerate single-direction predictors.
Minimum-movement filtering determines the effective difficulty
As shown in Figure˜2, performance is near chance without filtering or under weak filtering ($\tau\in\{0,0.002,0.004\}$), but improves sharply once $\tau$ reaches the mid range (notably $\tau\in\{0.006,0.008\}$), consistent with reduced label ambiguity on the ex-post filtered subsets. Aggregating across all deep configurations (Lite-CNN/ResNet18-P and *-cons, all input modes), the mean MCC rises from 0.002-0.002 at $\tau=0$ to $0.082$ at $\tau=0.006$ and peaks at $0.092$ at $\tau=0.008$, before dropping to $0.030$ at $\tau=0.010$. The drop at $\tau=0.010$ aligns with the reduced effective sample size and increased variance (see the sample-count/embargo summary in Table˜5, in Appendix˜D). We stress that $\tau$ defines offline evaluation regimes ex post, not an inference-time filtering rule; the filtering/splitting protocol is specified in Appendix˜D.
Fusion behavior once labels stabilize
Once $\tau\geq 0.006$, dual-view information becomes exploitable and both turns consistently positive for Lite-CNN family; however, under weak filtering (e.g., $\tau=0.004$) early fusion can exhibit negative transfer relative to the stronger single-view option (Table˜1). Across the two-view setting, late fusion is the dominant architectural gain over early fusion (representative numbers in Table˜1), while consistency regularization yields modest, $\tau$-dependent and non-monotonic effects.
Backbone differences
ResNet18-P exhibits a noisier regime: dual-view early fusion is generally weak relative to single-view modes at the same $\tau$, and late fusion partially recovers performance but with higher variance, especially under stricter filtering (see Figure˜2 and representative values in Table˜1).
LogReg as a diagnostic
LogReg can be competitive at certain thresholds but is non-monotonic across $\tau$ and input modes; we treat it primarily as a diagnostic indicator of partial linear separability after label stabilization.
Takeaways
Across backbones, (i) label stabilization via min_move is a prerequisite for learning in our setting; (ii) in the dual-view setting, late fusion is generally more reliable than early fusion on clean evaluation once labels stabilize; and (iii) consistency regularization can provide modest gains in some mid-noise regimes but is not reliably monotonic and requires careful tuning jointly with $\tau$. For full numerical commentary (without changing any conclusions or values), see Appendix˜G.
@@
Model (input) $\boldsymbol{\tau=0.004}$ $\boldsymbol{\tau=0.006}$ $\boldsymbol{\tau=0.008}$ Lite-CNN (both, early fusion) $020004\pm 0.016396 0.0$$088767\pm 0.042662$ $\mathbf{0.107206\pm 0.046632}$ Lite-CNN (indic) $\mathbf{0.107292\pm 0.022477}$ $0.079303\pm 0.032678$ $0.100279\pm 0.048322$ Lite-CNN (ohlcv) $0.003236\pm 0.023789$ $109095\pm 0.039023 0.0$$090707\pm 0.028893$ Lite-CNN-Cons (both, $\lambda=0$) $0.023262\pm 0.034602$ $0.127459\pm 0.035946$ $0.100430\pm 0.034439$ Lite-CNN-Cons (both, $\lambda=0.5$) $0.015634\pm 0.052001$ $\mathbf{0.129607\pm 0.026449}$ $0.089973\pm 0.067209$ Lite-CNN-Cons (both, $\lambda=1$) $0.014013\pm 0.034309$ $0.120755\pm 0.029468$ $0.103269\pm 0.031387$ ResNet18-P (both, early fusion) $012168\pm 0.030674 0.0$$17593\pm 0.044072 0.0$$042873\pm 0.061370$ ResNet18-P (indic) $055638\pm 0.047824 0.0$$082039\pm 0.039988$ $0.108023\pm 0.040145$ ResNet18-P (ohlcv) $005656\pm 0.059240 0.0$$018580\pm 0.054761$ $0.109101\pm 0.072830$ ResNet18-P-Cons (both, $\lambda=0.5$) $003653\pm 0.037078 0.0$$67112\pm 0.031388 0.0$$083423\pm 0.052584$
Choice of $\tau$ for robustness evaluation
All subsequent adversarial evaluations use $\tau=0.006$ as the default operating point. At smaller thresholds ($\tau\leq 0.004$), most configurations remain close to chance in MCC, making robustness comparisons less informative. At larger thresholds (e.g., $\tau=0.010$), the reduced effective sample size can amplify variance across random seeds. As a sensitivity check, we repeat the key robustness evaluations at $\tau=0.008$ and observe the same qualitative conclusions (see Appendix˜I). Thus, $\tau=0.006$ offers a practical balance between label stabilization and sufficient data for reliable robustness measurement.